A hospital switch board receives on an average 0.9 calls/minute. Find the probability that (i) no calls in a minute, (ii) 2 or more calls/minute



1
Expert's answer
2022-04-14T15:09:37-0400

Let's use Poisson's distribution "P(x=k)=\lambda^ke^{-\lambda}/k!,"

where "\lambda" is the mean

"1)\space P(x=0)=0.9^0e^{-0.9}/0!=\\
=1/e^{0.9}\approx0.4066=40.66\%\\
2)\space P(x\geq2)=1-P(x<2)=\\
=1-P(x=0)-P(x=1)=\\
=1-0.4066-0.9^1e^{-0.9}/1!\approx0.2275=22.75\%"